OpenStudy (anonymous):

TWO ELECTRIC BULBS WHOSE RESISTANCES ARE IN THE RATIO OF 1:2ARE CONNECTED IN PAALLEL TO ACONSTANT VOLTAGE SOURCE . THE POWER DISSIPATED IN THEM HAVE THE RATIO

OpenStudy (unklerhaukus):

http://openstudy.com/study#/groups/Physics

OpenStudy (unklerhaukus):



OpenStudy (marco26):

2:1, I guess

OpenStudy (marco26):

Since \[P=\frac{ V ^{2} }{ R }\]

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

\[I_1=2I_2\]

OpenStudy (unklerhaukus):

\[P=IV\]

OpenStudy (agent0smith):

You could sum the resistors first, to find... whatever it is you need to find, you didn't really say...?\[\Large \frac{ 1 }{ R _{T} }= \frac{ 1 }{ R _{1}}+\frac{ 1 }{ R _{2} }\] Where R1 and R2 are the R and 2R.